# Efficient Neural Common Neighbor for Temporal Graph Link Prediction

### v1

@@ 4.3 NCN-based Prediction Head
For source and destination nodes, we perform an element-wise product:
$X^{t}_{u,v}=emb_{u}^{t}\otimes emb_{v}^{t}.$ (2)
For multi-hop CN nodes, we aggregate their embeddings in each hop with sum pooling:
$NCN_{k}(u,v)=\mathop{\oplus}\limits_{w\in\text{CN}^{t}_{k}(u,v)}emb_{w}^{t}.$ (3)
@@
$repr(u,v)$ $=[X^{t}_{u,v}\ ||\ (\mathop{||}\limits_{k=0}^{K})NCN_{k}(u,v)].$ (4)
In the above, $\otimes$, $\oplus$, and ||||| | stand for element-wise product, element-wise summation, and concatenation of vectors, respectively. The pair-wise representation $repr(u,v)$ for nodes $u$ and $v$ is fed to a projection head to output the final link prediction.
@@ 5 Efficiency and Effectiveness of TNCN
In this section, we explore the two principal benefits of TNCN: efficiency and effectiveness. These advantages are demonstrated through an analysis of two core components within the framework for temporal graph link prediction: graph representation learning and link prediction methods. We categorize graph representation learning modules into two types: memory-based and $k$-hop-subgraph-based, according to the temporal scope of evolved events. Memory-based modules exhibit superior time complexity while maintaining good expressiveness in some situations, striking a balance between efficiency and performance. Furthermore, we highlight the deficiencies of existing link prediction methods on temporal graph learning and introduce the extended common neighbor approach. This method serves as a complementary addition for learning pair-wise representations while eliminating the necessity for message passing on entire graphs. Both our graph representation learning and link prediction techniques are designed with a unified optimization objective: to avoid message passing on entire subgraphs in favor of non-repetitive operations, culminating in a cohesive solution that is both efficient and effective.
5.1 Graph Representation Learning
Graph representation learning aims to develop an embedding function, denoted as $Emb$, which learns an embedding for each node encoding its structural and feature information within the graph. Specifically, given a new event represented as $(u,v,t)$, the function $Emb$ leverages prior events to generate meaningful embeddings. Approaches in this domain diverge in their handling of temporal dynamics; some opt to maintain a dynamic embedding for each node that is incrementally updated with each new event, while others choose to recalculate node embeddings by considering the entire historical context of events, thereby providing a more comprehensive reflection of past interactions. We classify these methodologies into two distinct types based on their operational mechanisms.
@@ 5.1 Graph Representation Learning
Memory-based approach. Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is referred to as a memory-based approach.
$Emb(u,t)=f_{\textit{emb}}(Mem(u,t^{\prime})),~{}$ $Mem(u,t)=f_{\textit{mem}}(Mem(u,t^{\prime}),Mem(v,t^{\prime}),e_{u,v}^{t},t-t^{\prime}),$ (5)
$Emb(v,t)=f_{\textit{emb}}(Mem(v,t^{\prime})),~{}$ $Mem(v,t)=f_{\textit{mem}}(Mem(v,t^{\prime}),Mem(u,t^{\prime}),e_{v,u}^{t},t-t^{\prime}),$
where $f_{\textit{emb}}$ and $f_{\textit{mem}}$ are two learnable functions.
@@
$k$-hop-subgraph-based approach. Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is defined as a subgraph-based approach.
@@
Effectiveness.
The analysis begins by assessing the effectiveness of the two paradigms. To do so, we first introduce the concept of $k$-hop event (LOV´ASZ et al., 1993).
@@
($k$-hop event & monotone $k$-hop event) A $k$-hop event is a sequence of consecutive edges $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$ connecting the initial node $u_{0}$ to the final node $u_{k}$. For example, $\{(u,x,t^{\prime}),(x,v,t)\}$ is a $2$-hop event. In the case where $k=1$, the $k$-hop event reduces to a single interaction $(u,v,t)$. A monotone $k$-hop event is a $k$-hop event in which the sequence of timestamps $\{t_{u_{i},u_{i+1}}\mid i\in\{0,\ldots,k-1\},k\geq 1\}$ is strictly monotonically increasing.
@@
From Theorem 5.4, we can conclude that 1) memory-based approaches have superior expressiveness in encoding $k$-hop events compared to 1-hop-subgraph-based approaches, 2) memory-based approaches have superior expressiveness in encoding monotone $k$-hop events than $k^{\prime}$-hop-subgraph-based approaches when $k^{\prime}<k$, and 3) $k$-hop-subgraph-based approaches are not less expressive than memory-based approaches when $k$ is large enough.
@@
Efficiency.
We then turn our attention to the efficiency of the two approaches. A pivotal factor is the frequency with which individual events are incorporated into computations. In memory-based approaches, each event is utilized a single time for learning, immediately following its associated prediction. Conversely, in the $k$-hop-subgraph-based method, an event may be employed multiple times, as it is revisited in different nodes’ temporal neighborhood and repeated been processed within each subgraph’s encoding (such as message passing) process. This discrepancy leads to divergent cumulative frequencies of event utilization throughout the learning process and results in the huge efficiency advantage of memory-based methods. We formalize this observation in the following:
@@
Following Theorem 5.5, it becomes evident that the computational overhead incurred by a memory-based method is significantly lower than that of a $k$-hop-subgraph-based method, particularly as the value of $k$ increases. These results highlight the advantages of memory-based methods in mitigating the computational efficiency challenges associated with large-scale temporal graphs.
5.2 Link Prediction Technique
With the node embeddings obtained from the graph representation learning step, link prediction techniques involve aggregating the node embeddings in some way into link representations for link prediction. Most previous methods simply concatenate the source and destination node embeddings as the link representation $Emb(u,t)\mathop{||}Emb(v,t)$, losing a great amount of pairwise features (as illustrated in Figure 1). In order to use labeling trick, an alternative approach involves extracting a separate subgraph for each link to predict, which leads to good results but also significantly increases the complexity. TNCN, on the contrary, utilizes neural common neighbors as a decoupled and flexible solution. This approach can be seamlessly integrated into existing memory-based methods with minimal computational overhead, while retaining important structural information for link prediction.
Effectiveness.
In the following, we first demonstrate the effectiveness of TNCN by showing that it can capture three important pairwise features commonly used as effective link prediction heuristics, namely Common Neighbors (CN), Resource Allocation (RA), and Adamic-Adar (AA) (Newman, 2001; Adamic and Adar, 2003; Zhou et al., 2009), borrowing from NCN (Wang et al., 2023).
Theorem 5.6.
TNCN is strictly more expressive than CN, RA, and AA.
Essentially, TNCN uses node memory to substitute the constant or degree-based values in the three features, and the memory update scheme is sufficient to learn such values. In comparison, an approach that merely concatenates node-wise representations proves inadequate in capturing these heuristics.
Efficiency.
We then address the efficiency of TNCN. The computation of TNCN can be divided into two primary components: the generation of neighbor nodes’ embeddings and the execution of common neighbor lookups. Concerning the former, the memory-based approach intrinsically tracks all node embeddings, thus obviating the need for re-computation. As for the latter, we implement a fast common neighbor search algorithm by leveraging a sparse matrix structure that supports batch operations. Collectively, these factors contribute to the minimal additional overhead of TNCN compared to memory-based methods.
@@ 6.1 Experimental Settings
Datasets
We evaluate our model on five large-scale real-world datasets for temporal link prediction from the Temporal Graph Benchmark (Huang et al., 2023). These datasets span several distinct fields: co-editing network on Wikipedia, Amazon product review network, cryptocurrency transactions, directed reply network of Reddit, and crowdsourced international flight network. They vary in scales and time spans. Additional details about the datasets are provided in Appendix A. We set the evaluation metric as Mean Reciprocal Rank (MRR) consistent with the TGB official leaderboard.
Baselines
We systematically evaluate our proposed model against a diverse set of baselines known for their strong capacity to represent temporal graph dynamics. These include: a heuristic algorithm Edgebank (Yu et al., 2023), memory-based models DyRep (Trivedi et al., 2019) and TGN (Rossi et al., 2020) that obviate the need for frequent temporal subgraph sampling, and GraphMixer (Cong et al., 2023) which primarily employs an MLP-mixer. Additionally, we include various graph-based models such as CAWN (Wang et al., 2021c), NAT (Luo and Li, 2022), DyGFormer (Yu et al., 2023), TGAT (Xu et al., 2020), and TCL (Wang et al., 2021a), which learn from neighborhood structure information. To reconcile the trade-off between model efficacy and computational overhead, we evaluate two TNCN variants: the “TNCN-standard” (TNCN(S)), which incorporates up to 1-hop common neighbor information, and the more complex “TNCN-large” (TNCN(L)), which uses up to 2-hop common neighbors.
@@ 6.2 Experimental Results
Table 1: Test Performance of different models under MRR metric. The top three are emphasized by red, blue and black fonts. ‘NA’ denotes scenarios where a specific method was either not applied to the dataset or was unable to complete the validation and testing phases within a reasonable timeframe.
Model Wiki Review Coin Comment Flight
TGN 0.528 0.387 0.737 0.622 0.705
DyGformer 0.798 0.224 0.752 0.670 NA
NAT 0.749 0.341 NA NA NA
CAWN 0.711 0.193 NA NA NA
Graphmixer 0.118 0.521 NA NA NA
TGAT 0.141 0.355 NA NA NA
TCL 0.207 0.193 NA NA NA
DyRep 0.050 0.220 0.452 0.289 0.556
Edgebank(tw) 0.571 0.025 0.580 0.149 0.387
Edgebank(un) 0.495 0.023 0.359 0.129 0.167
TNCN-standard 0.720 0.298 0.739 0.692 0.803
TNCN-large 0.717 0.317 0.764 0.662 NA
Reply to Q1: TNCN possesses remarkable performance.
We conducted comprehensive evaluations of prevailing methods on the TGB. The main results are summarized in Table 1. It is evident from the table that TNCN attains new SOTA performance on three out of five datasets. Additionally, TNCN demonstrates competitive results on the remaining two datasets. TNCN almost consistently surpasses memory-based models such as TGN and DyRep, which can be attributed to its utilization of supplementary structural information. Even in comparison to powerful graph-based models, including NAT and DyGFormer, TNCN still matches or exceeds their performance, underscoring the effectiveness of its integration of node memory and pair-wise features. The only dataset where TNCN exhibits relatively poorer performance is the Review dataset. This can be ascribed to the dataset’s predominant reliance on feature embedding and its high "surprise" metric, wherein prior events have a diminished correlation with subsequent events, potentially reducing the impact of CNs.
Reply to Q2: TNCN shows great scalability on large datasets.
@@
Figure 3: Time Consumption of Memory and Graph-based Method on Wiki and Review Datasets
To evaluate computational efficiency, we collected the time consumption when applying all baselines to the Wiki and Review datasets, as depicted in Figure 3. Compared with memory-based methods, TNCN exhibits a comparable order of magnitude in terms of time consumption. However, when benchmarked against graph-based models, TNCN demonstrates a substantial acceleration, achieving approximately 1.9 to 4.7 times speedup during the training phase and a 2.1 to 6.4 times increase in inference speed. Notably, the scalability concerns become even more evident as the size of the dataset expands; several graph-based models cannot complete the validation and testing processes within a reasonable time budget. The primary factors contributing to TNCN’s efficiency are the synergistic, time-efficient design of its two core components and the implementation of the Efficient CN Extractor that facilitates batch operations through parallel processing.
Reply to Q3: Extended CN brings improvements.
Table 2: Test performance of TGN and TNCN with different ranges of common neighbors

### v2

@@ 3.3 NCN-based Prediction Head
We finally construct our NCN-based representation as follows. For source and destination nodes, we perform an element-wise product. For multi-hop CN nodes, we aggregate their embeddings in each hop with sum pooling.
$X^{t}_{u,v}=emb_{u}^{t}\otimes emb_{v}^{t},\quad NCN_{(i,j)}(u,v)=\mathop{\oplus}\limits_{w\in\text{CN}^{t}_{(i,j)}(u,v)}emb_{w}^{t}.$ (3)
@@
$repr(u,v)$ $=[X^{t}_{u,v}\ ||\ (\mathop{||}\limits_{i,j}^{K})NCN_{(i,j)}(u,v)].$ (4)
We have used $\otimes$, $\oplus$, and $||$ to denote element-wise product, element-wise summation, and concatenation of vectors, respectively. The pair-wise representation $repr(u,v)$ for nodes $u$ and $v$ will be fed to a projection head to output the final link prediction.
@@ 4 Efficiency and Effectiveness of TNCN
In this section, we explore the two principal benefits of TNCN: efficiency and effectiveness. These advantages are demonstrated through an analysis of two core components within the framework for temporal graph link prediction: graph representation learning and link prediction methods.
Temporal graph representation learning aims to develop an embedding function, denoted as $Emb$, which learns an embedding for each node encoding its structural and feature information within the graph. Specifically, given a new event represented as $(u,v,t)$, the function $Emb$ leverages prior events to generate meaningful embeddings. We first categorize graph representation learning approaches into two types: memory-based and $k$-hop-subgraph-based, according to their temporal scope of evolved events.
@@
Memory-based approach. Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is referred to as a memory-based approach, which opts to maintain a dynamic, incrementally updated embedding for each node.
$Emb(u,t)=f_{\textit{emb}}(Mem(u,t^{\prime})),~~~~Emb(v,t)=f_{\textit{emb}}(Mem(v,t^{\prime})),$ (5)
where the $Mem$ can be obtained from the pipeline in Definition 2.4, and $f_{\textit{emb}}$ is a learnable function.
@@
$k$-hop-subgraph-based approach. Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is defined as a subgraph-based approach, which chooses to recalculate node embeddings by considering the entire historical events.
@@
4.1 Effectiveness
The analysis begins by assessing the effectiveness of the two paradigms. To do so, we first introduce the concept of $k$-hop event [LOVASZ1993RandomWO].
@@ 4.1 Effectiveness
$k$-hop event & monotone $k$-hop event. A $k$-hop event is a sequence of consecutive edges $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$ connecting the initial node $u_{0}$ to the final node $u_{k}$. For example, $\{(u,x,t^{\prime}),(x,v,t)\}$ is a $2$-hop event. In the case where $k=1$, the $k$-hop event reduces to a single interaction $(u,v,t)$. A monotone $k$-hop event is a $k$-hop event in which the sequence of timestamps $\{t_{u_{i},u_{i+1}}\mid i\in\{0,\ldots,k-1\},k\geq 1\}$ is strictly monotonically increasing.
@@
The proof is in Appendix I. From Theorem 4.4, we can conclude that 1) memory-based approaches have superior expressiveness in encoding $k$-hop events compared to 1-hop-subgraph-based approaches, and 2) memory-based approaches have superior expressiveness in encoding monotone $k$-hop events than $k^{\prime}$-hop-subgraph-based approaches when $k^{\prime}<k$.
@@
Corollary 4.5.
If we use up to $k$ hop neighborhood information of central node $u$, then TNCN can capture at least $(k+1)$-hop subgraph information around $u$.
This is because TNCN with memory-based backbone can obtain additional 1-hop information regardless of the time monotony, i.e. arbitrary central node can interact with its neighbor when the edge between them exists. This can extend TNCN’s capability for free.
We also demonstrate the effectiveness of TNCN in the following theorem 4.6 with the proof in Appendix I. The first part shows that it can capture three important pairwise features commonly used as effective link prediction heuristics, namely Common Neighbors (CN), Resource Allocation (RA), and Adamic-Adar (AA) [newman2001clustering, adamic2003friends, zhou2009predicting]. The experimental results in Appendix D further validate this claim. In the second part we reveal that TNCN is strictly more expressive than some traditional temporal graph networks including Jodie [kumar2019predicting], DyRep [trivedi2019dyrep], TGN [rossi2020temporal] and TGAT [xu2020inductive] under the same condition, which are widely used as baselines.
Theorem 4.6.
(Expressivity of TNCN)
  1. 1.
    TNCN is strictly more expressive than CN, RA, and AA.
  2. 2.
    TNCN is strictly more expressive than Jodie with the same dimension of time encoding, DyRep with the same aggregation function, TGAT with the same attention layers and neighbors, and TGN under identical condition for all module choices.
From this theorem we can find that TNCN extends the previous generic memory-based framework of temporal graph networks with explicitly adopting the neural embeddings of common neighbors. This method can serve as a complementary addition for learning pair-wise representations.
4.2 Efficiency
We then turn our attention to the efficiency of the two approaches. A pivotal factor is the frequency with which individual events are incorporated into computations. In memory-based approaches, each event is utilized a single time for learning, immediately following its associated prediction. Conversely, in the $k$-hop-subgraph-based method, an event may be employed multiple times, as it is revisited in different nodes’ temporal neighborhood and repeated been processed within each subgraph’s encoding (such as message passing) process. This discrepancy leads to divergent cumulative frequencies of event utilization throughout the learning process, resulting in the huge efficiency advantage of memory-based methods. We formalize this observation as:
@@ 4.2 Efficiency
The proof is attached in Appendix I with part of the proof based on a classic conclusion from DECAEN1998245 in the graph theory. Following Theorem 4.7, it becomes evident that the computational overhead incurred by a memory-based method is significantly lower than that of a subgraph-based method, particularly as $k$ increases. These results highlight the advantages of memory-based methods in mitigating the computational efficiency challenges associated with large-scale temporal graphs.
So based on the memory backbone, our TNCN furthur utilizes an efficient CN extractor, eliminating the necessity for message passing on entire graphs. Consequently it achieves a unified optimization objective: to avoid message passing on entire subgraphs in favor of non-repetitive operations.
To summarize, TNCN introduces the extended common neighbor approach with the efficient CN extractor. This method serves as a complementarity for learning pair-wise representations, while culminating in a cohesive solution that is both efficient and effective.
5 Related Work
Memory-based Temporal Graph Representation Learning. Memory-based models learn node memory using continuous events with non-decreasing timestamps. Researchers have proposed several memory-based methods including JODIE [kumar2019predicting], DyRep [trivedi2019dyrep] and TGN [rossi2020temporal]. These methods are superior in higher efficiency, while lacking in capturing structural information.
Graph-based Temporal Graph Representation Learning. Subsequent works have incorporated the neighborhood structure into temporal graph learning. CNE-N [cheng2024co] employs a hash table to map interactions and computes co-neighbor encodings to predict future links. Unlike our TNCN which directly utilizes CN embeddings in prediction, CNE-N simply counts neighbor numbers. Additionally, while CNE-N manages recent interactions through hash tables, TNCN adopts a monotonic storage scheme. NAT [luo2022neighborhood] similarly builds a multi-hop node dictionary to compress neighbors, but it may suffer from hash collisions. DyGFormer [yu2023towards] encodes one-hop neighbors and their co-occurrences, then uses a Transformer for predictions, requiring repeated neighborhood sampling and computations. It only models 1-hop neighbor frequency and does not use CN embeddings. In contrast, TNCN supports multi-hop neighbors and includes their embeddings, enabling richer representations.
We also include detailed introduction to more related works in Appendix B.
@@ 6.1 Experimental Settings
Datasets. We evaluate our model on five large-scale real-world datasets for temporal link prediction from the Temporal Graph Benchmark [huang2023temporal]. These datasets span several distinct fields: co-editing network on Wikipedia, Amazon product review network, cryptocurrency transactions, directed reply network of Reddit, and crowdsourced international flight network. They vary in scales and time spans. Additional details about the datasets are provided in Appendix A. We set the evaluation metric as Mean Reciprocal Rank (MRR) consistent with the TGB official leaderboard.
Baselines. We systematically evaluate our proposed model TNCN against a diverse set of baselines: a heuristic algorithm Edgebank [yu2023towards], memory-based models Jodie [kumar2019predicting], DyRep [trivedi2019dyrep] and TGN [rossi2020temporal] that obviate the need for frequent temporal subgraph sampling, and GraphMixer [cong2023we] which employs an MLP-mixer. We also include various graph-based models such as CAWN [wang2021inductive], TGAT [xu2020inductive], TCL [wang2021tcl], NAT [luo2022neighborhood], DyGFormer [yu2023towards], CNE-N [cheng2024co] and TPNet [lu2024improvingtemporallinkprediction], which learn from neighborhood structure information.
Here we evaluate our TNCN under two similar but different settings, the official setting (“official”) and the new setting (“ns”). “*-official” strictly complies to the official setting of TGB evaluation policy, using both streaming setting and lag-one scheme for both memory update and neighborhood awareness. Streaming setting means the information of the validation and test sets can only be employed for updating the memory without any back propagation. Lag-one scheme implies that the model can access only the information from before the current batch for predictions; in other words, the latest usable batch is the previous one. This applies to not only the memory, but also the neighborhood awareness. “*-ns” obeys the streaming setting but considers the interactions within the same batch before the current prediction time. This allows the model to utilize more recent neighborhood information, potentially giving it unfair advantages in datasets where recent interactions are crucial. Methods “*-official” use the former setting while others report the latter. For a fair evaluation and comparison, here we display the performance of our TNCN under both settings.
Table 1: Test Performance of different models under MRR metric. The top three are emphasized by red, blue and bold fonts. ‘-’ denotes scenarios where a specific method was either not applied to the dataset or was unable to complete the validation and testing phases within a reasonable timeframe.
Model Wiki Review Coin Comment Flight
JODIE 0.631 $\pm$ 1.69 0.414 $\pm$ 0.15 - - -
DyRep 0.519 $\pm$ 1.95 0.401 $\pm$ 0.59 0.452 $\pm$ 4.60 0.289 $\pm$ 3.30 0.556 $\pm$ 1.40
TGAT 0.599 $\pm$ 1.63 0.196 $\pm$ 0.23 0.609 $\pm$ 0.57 0.562 $\pm$ 2.11 -
TGN-official 0.528 $\pm$ 0.06 0.387 $\pm$ 0.02 0.737 $\pm$ 0.03 0.622 $\pm$ 0.02 0.705 $\pm$ 0.02
TGN-ns 0.689 $\pm$ 0.53 0.375 $\pm$ 0.23 0.586 $\pm$ 3.70 0.379 $\pm$ 2.10 -
CAWN 0.730 $\pm$ 0.60 0.193 $\pm$ 0.10 - - -
EdgeBank(tw) 0.633 0.029 0.574 0.149 0.387
EdgeBank(un) 0.525 0.023 0.359 0.129 0.167
TCL 0.781 $\pm$ 0.20 0.165 $\pm$ 1.85 0.687 $\pm$ 0.30 0.701 $\pm$ 0.83 -
GraphMixer 0.598 $\pm$ 0.39 0.369 $\pm$ 1.50 0.756 $\pm$ 0.27 0.762 $\pm$ 0.17 -
NAT 0.749 $\pm$ 1.00 0.341 $\pm$ 2.00 - - -
DyGFormer 0.798 $\pm$ 0.42 0.224 $\pm$ 1.52 0.752 $\pm$ 0.38 0.670 $\pm$ 0.14 -
CNE-N 0.802 $\pm$ 0.20 0.261 $\pm$ 0.25 0.772 $\pm$ 0.21 0.790 $\pm$ 0.14 -
TPNet 0.827 $\pm$ 0.01 - 0.832 $\pm$ 0.01 0.825 $\pm$ 0.06 0.884 $\pm$ 0.01
TNCN-official 0.724 $\pm$ 0.01 0.419 $\pm$ 0.09 0.770 $\pm$ 0.06 0.727 $\pm$ 0.12 0.817 $\pm$ 0.04
TNCN-ns 0.803 $\pm$ 0.01 0.427 $\pm$ 0.06 0.771 $\pm$ 0.04 0.705 $\pm$ 0.12 0.831 $\pm$ 0.03
@@ 6.2 Experimental Results
Reply to Q1: TNCN possesses remarkable performance. We conducted comprehensive evaluations of prevailing methods on TGB. The main results are summarized in Table 1. It is evident from the table that TNCN attains new SOTA performance on Review dataset. Additionally, TNCN demonstrates competitive results on the remaining datasets, ranking 2nd or/and 3rd on Wiki, Coin and Flight. TNCN almost consistently surpasses both memory-based models such as TGN and DyRep, and graph-based methods like DyGFormer and NAT. The dataset where TNCN still exhibits a large gap from the top-three baselines is Comment. This may be ascribed to the high “surprise index” of this non-bipartite dataset, wherein prior events have a diminished correlation with subsequent events, potentially reducing the impact of CNs. (More details about Surprise index are in Appendix A.)
To obtain a more comprehensive evaluation, we have also conducted experiments on the previous small and medium datasets under both transductive and inductive settings. In these experiments, TNCN achieves 6 new SOTA out of 7 datasets in transductive setting and 3 in inductive setting, further facilitating its strong performance. The overall performance and some additional results (such as TGN with heuristics, etc.) can be referred to Appendix D.
Reply to Q2: TNCN shows great scalability on large datasets. To evaluate computational efficiency, we collected the time consumption on Wiki and Review datasets, as depicted in Figure 3. Compared with memory-based methods, TNCN exhibits a comparable order of magnitude in terms of time consumption. However, when benchmarked against graph-based models, TNCN demonstrates a substantial acceleration, achieving approximately 2.5 to 5.9 times speedup during the training phase and a 1.8 to 30.3 times increase in inference speed. In the Table  13 we provide a full comparison of the time consumption for different models over the five TGB datasets. Notably, the scalability concerns become even more evident as the size of the dataset expands; several graph-based models cannot complete the validation and testing processes within a reasonable time budget. The primary factors contributing to TNCN’s efficiency are the synergistic, time-efficient design of its two core components and the implementation of the Efficient CN Extractor that facilitates batch operations through parallel processing. For detailed statistics about TNCN and NAT, please refer to Appendix E.1.
@@
Figure 3: Time Consumption of Memory and Graph-based Method on Wiki and Review Datasets.
6.3 Ablation Study
Reply to Q3: Extended CN brings improvements. To elucidate the benefits of extended CN, we conducted an ablation study under official setting on the hop range of common neighbors. The results are shown in Table 2. Here we use notation "$k$-hop CN" to simply denote the CNs up to $(k,k)$-hop. The conventional NCN method considers only (1,1)-hop CN. However, this approach may not be universally applicable across all temporal networks. For instance, bipartite graphs lack such (1,1)-hop CN in their structure, necessitating the consideration of 2-hop CN. Additionally, memory-based methods may omit a notable aspect: they generally find it difficult to quantify the frequency of interactions between a given pair of nodes, which brings the need for 0-hop neighborhood.
[…]
Table 2: Test performance of TGN and TNCN with different ranges of common neighbors.